350 3000.964 0.925 0.899 0.805 0.755 250 0.964 0.926 0.900 O.81 0.76 нсо, H,citrate-, CH,NH, (CHINHİ K, cr, Br-, I-, CN-,NOİ, NOİ 0.964 0.924 0.898 0.80 0.75 Accounting for ionic strength, what is the solubility (moles/L) of Pb(BrOs)2 in water that also contains 0.0500 M NaCI and 0.0500 M KBr? NaCI and KBr are both fully soluble. The appropriate activity coefficient (v) for Pb* is 0.37; you need to find the activity coefficient for BrOs in the table below. The Ksp for Pb(BrO3)h is 2.1 x 106.
Answer #1

ionic strength of solution (u) = 1/2Ci.zi^2

u = 1/2(0.05 x 1^2 + 0.05 x 1^2 + 0.05 x 1^2 + 0.05 x 1^2)

   = 0.1

Using extended debye-huckel equation

Activity coefficient [Y] for BrO3-

= inv. log[-0.51 x 1^1 x sq.rt.(0.1)/(1+3.3 x 0.154 x sq.rt.(0.1))]

= 0.726

Activity coefficient [Y] for Pb2+ = 0.370

let x amount of Pb(BrO3)2 is in solution

then,

Ksp = [Pb2+][BrO3-]

2.1 x 10^-6 = (x)(0.370)(2x)^2.(0.726)^2

x = 0.014 M

molar solubility = 0.014 M